OpenStudy (loser66):

Statistics question. Please, help Estimate \(\mu\) \(X_i\)~ \(N(\mu, \sigma^2)\) \(\hat\mu=\dfrac{X_1+\cdots+X_n}{n}\)

OpenStudy (unklerhaukus):

What is \(N\)

OpenStudy (loser66):

In class, my Prof did \(E(\hat\mu)= \mu\) I understood it. \(Var(\hat\mu)=\dfrac{\sigma^2}{n}\). I understood it also And then \(B_\mu \hat\mu =\mu-\mu =0\) I got it also Then \(MSE_\mu \hat\mu= Var(\hat\mu)+B_\mu \hat\mu = \dfrac{\sigma^2}{n}\)

OpenStudy (loser66):

And said, as n goes to infinitive, MSE goes to 0. and done

OpenStudy (loser66):

I don't get why he stopped there. The question is about estimate \(\mu\), I didn't see it at the end.

OpenStudy (loser66):

@Zarkon

OpenStudy (math&ing001):

If I understand well, the MSE goes to 0 when we have an infinite amount of samples (n goes to infinitive). And, in this case, we can say that the estimator û predicts observations of the parameter μ with perfect accuracy (i.e we can estimate μ by û).

OpenStudy (loser66):

Thanks for explanation.

OpenStudy (math&ing001):

Welcome !

OpenStudy (zarkon):

In your book...loot at definition 9.2 and theorem 9.1 on page 450

OpenStudy (loser66):

It means we use this method to consider whether \(\hat\mu\) is consistence or not, right?